# Gromov-Hausdorff convergence of metric spaces of UCP maps

### v1

@@ 4.3. The $d$-dimensional torus
where $\mathcal{H}$ and $D$ are given by (4.3). Note that the truncations for the torus are different from the truncation for the unit circle discussed above. We recall the following from [8].
@@
$\alpha_{\theta}(T)=\left(t_{k-l}e^{i(k-l)\cdot\theta}\right)_{k,l\in\overline{\mathrm{B}}_{N}^{\mathbb{Z}}},$ (4.5)
where $\theta=(\theta_{1},\theta_{2},\cdots,\theta_{d})$ such that $(e^{i\theta_{1}},e^{i\theta_{2}},\cdots,e^{i\theta_{d}})\in\Omega$, the vector $|\psi\rangle$ is given by $|\psi\rangle=\sum_{n\in\overline{\mathrm{B}}_{N}^{\mathbb{Z}}}e_{n}$, and $\mathcal{N}_{\mathrm{B}}(N):=\#\overline{\mathrm{B}}_{N}\cap\mathbb{Z}^{d}$ is the number of $\mathbb{Z}^{d}$-lattice points in the closed ball of radius $N$. By Lemma 2.3 and Lemma 2.6, $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$ have BW-property.
It is shown in the proof of [8, Theorem 3.10] that the pair $(R_{N},S_{N})$ form a $C^{1}$-approximate order isomorphism for the spectral triples $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$. By [9, Theorems 3.9 and 3.11], $R_{N}$ and $S_{N}$ are ucp maps, and hence $(R_{N},S_{N})$ form a $C^{1}$-approximate complete order isomorphism. Therefore, applying Theorem 1, we have the following convergnce result.

### v2

@@ 4.3.1. Spherical truncation
where $\mathcal{H}$ and $D$ are given by (4.1). We call the above truncation (4.2) the spherical truncation of the (4.1). Note that the truncations for the torus are different from the truncation for the unit circle discussed above. We recall the following from [11].
@@
$\alpha_{\theta}(T)=\left(t_{k-l}e^{i(k-l)\cdot\theta}\right)_{k,l\in\overline{\mathrm{B}}_{N}^{\mathbb{Z}}},$
where $\theta=(\theta_{1},\theta_{2},\cdots,\theta_{d})$ such that $(e^{i\theta_{1}},e^{i\theta_{2}},\cdots,e^{i\theta_{d}})\in\Omega$, the vector $|\psi\rangle$ is given by $|\psi\rangle=\sum_{n\in\overline{\mathrm{B}}_{N}^{\mathbb{Z}}}e_{n}$, and $\mathcal{N}_{\mathrm{B}}(N):=|\overline{\mathrm{B}}_{N}\cap\mathbb{Z}^{d}|$ is the number of $\mathbb{Z}^{d}$-lattice points in the closed ball of radius $N$. By Lemma 2.3 and Lemma 2.6, $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$ have the BW-property.
It is shown in the proof of [11, Theorem 3.10] that the pair $(R_{N},S_{N})$ form a $C^{1}$-approximate order isomorphism for the spectral triples $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$. By [12, Theorems 3.9 and 3.11], $R_{N}$ and $S_{N}$ are ucp maps, and hence the pair $(R_{N},S_{N})$ form a $C^{1}$-approximate complete order isomorphism. Therefore, applying Theorem 1, we have the following convergnce result.
@@
4.3.2. Polyhedral truncation
This section is inspired by [17].
Definition 4.4.
Let $r,d\in\mathbb{N}$. A polyhedron $\widetilde{\Delta}(r)$ in $\mathbb{R}^{d}$ is the convex hull of a finite set $W=\{\bm{a}_{1},\ldots,\bm{a}_{r}\}\subset\mathbb{Z}^{d}$, where $W$ is minimal, meaning that it coincides with the set of vertices of $\widetilde{\Delta}(r)$. The restriction $\Delta(r)=\widetilde{\Delta}(r)\cap\mathbb{Z}^{d}$ is called a summability polyhedron if $\widetilde{\Delta}(r)$ contains the origin strictly in its interior. For any $N\in\mathbb{N}$, we denote by $\widetilde{\Delta}_{N}(r)$ the convex hull of the set $\{N\bm{a}_{1},\ldots,N\bm{a}_{r}\}$ and define $\Delta_{N}(r)=\widetilde{\Delta}_{N}(r)\cap\mathbb{Z}^{d}$. When the polyhedron is fixed, that is, $r$ and $W$ are fixed, we simply write $\Delta$ and $\Delta_{N}$ instead of $\Delta(r)$ and $\Delta_{N}(r)$, respectively. We call $\Delta_{N}$ the $N$-dilation of $\Delta$.
Let $\Delta$ be a summability polyhedron and let $\Delta_{N}$ be the $N$-dilation of $\Delta$ for an integer $N\geq 1$. Let $Q_{N}$ be the orthogonal projection to the subspace of $\mathcal{H}$ spanned by the eigenspinors $e_{\lambda}$ with $\lambda\in\Delta_{N}$, i.e., $Q_{N}\mathcal{H}=\mathrm{span}\{e_{\lambda}\,:\lambda\in\Delta_{N}\}\otimes V$, with $e_{n}(x):=e^{in\cdot x}$, for all $x\in\Omega$.
Let $\mathcal{T}^{d}_{N}(PH)$ (PH stands for polyhedron) be the operator system of $d$-dimensional polyhedral Toeplitz matrices $T=(t_{k-l})_{k,l\in\Delta_{N}}$. This is a $|\Delta_{N}|\times|\Delta_{N}|$ matrix with entries indexed by the elements of $\Delta_{N}$. The term $t_{k-l}$ represents the inner product $\langle e_{k},Te_{l}\rangle$.
The spectral projection $Q_{N}$ defines the operator system spectral triple:
$\left(\mathcal{E}_{N}:=\mathcal{T}^{d}_{N}(PH),\mathcal{H}_{N}:=Q_{N}\mathcal{H},D_{N}:=Q_{N}D|_{\mathcal{H}_{N}}\right),$ (4.3)
where $\mathcal{H}$ and $D$ are given by (4.1). We call the above truncation (4.3) a polyhedral truncation of (4.1). Define
$\rho_{N}:$ $~{}\mathcal{E}\to\mathcal{E}_{N}$
$f\mapsto Q_{N}fQ_{N},$
$\sigma_{N}:$ $~{}\mathcal{E}_{N}\to\mathcal{E}$
$T\mapsto\frac{1}{|\Delta_{N}|}\mathrm{Tr}\left(|\psi\rangle\langle\psi|\alpha(T)\right).$
Here, $\alpha$ is the $\Omega$-action on $\mathcal{E}_{N}$ given by
$\alpha_{\theta}(T)=\left(t_{k-l}e^{i(k-l)\cdot\theta}\right)_{k,l\in\Delta_{N}},$
where $\theta=(\theta_{1},\theta_{2},\cdots,\theta_{d})$ is such that $(e^{i\theta_{1}},e^{i\theta_{2}},\cdots,e^{i\theta_{d}})\in\Omega$ and the vector $|\psi\rangle$ is given by $|\psi\rangle=\sum_{n\in\Delta_{N}}e_{n}$. By Lemma 2.3 and Lemma 2.6, $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$ have the BW-property.
Using the approach of [11], we show below that the pair $(\rho_{N},\sigma_{N})$ forms a $C^{1}$-approximate order isomorphism for the spectral triples $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$.
Lemma 4.5.
The maps $\rho_{N}$ and $\sigma_{N}$ are unital, completely positive and $\mathrm{C}^{1}$-contractive.
Proof.
The proof is similar to [11, Lemma 3.1 and Lemma 3.2], along with an application of [12, Theorems 3.9 and 3.11] as $C^{\infty}(\Omega)$ is commutative. ∎
Proposition 4.6.
The pair $(\rho_{N},\sigma_{N})$ forms a $C^{1}$-approximate order isomorphism for the spectral triples $(\mathcal{E}_{N},\mathcal{H}_{N},D_{N})$ and $(\mathcal{E},\mathcal{H},D)$ given by (4.3) and (4.1), respectively.
Proof.
By a computation similar to [11, Lemma 3.3], we have
$\sigma_{N}\circ\rho_{N}(f)(x)=\sum_{n\in\Delta_{N}-\Delta_{N}}\frac{|\Delta_{N}\cap(\Delta_{N}+n)|}{|\Delta_{N}|}\widehat{f}(n)e_{n}(x)$
$\rho_{N}\circ\sigma_{N}(T)=\left(\frac{|\Delta_{N}\cap(\Delta_{N}+m-n)|}{|\Delta_{N}|}t_{m-n}\right)_{m,n\in\Delta_{N}}$
For each $N\in\mathbb{N}$, define $K_{2N}$ on $\Omega$ by
$K_{2N}(x)=$ $\sum_{n,m\in\Delta_{N}}\frac{1}{|\Delta_{N}|}e_{n-m}(x)$
$=$ $\sum_{n\in\Delta_{N}-\Delta_{N}}\frac{|\Delta_{N}\cap(\Delta_{N}+n)|}{|\Delta_{N}|}e_{n}(x).$
Then it is easy to check that
$\sigma_{N}\circ\rho_{N}(f)(x)=(K_{2N}*f)(x).$
A careful examination of the proof of [11, Theorem 3.10] suggests that, to demonstrate that $(\rho_{N},\sigma_{N})$ form a $C^{1}$-approximate order isomorphism, it suffices to show that the family $\{K_{2N}\}_{N\in\mathbb{N}}$ serves as a good kernel. In [17], $K_{2N}$ is called the polyhedron Fejér kernel. From [17, Theorems 2, 3], it follows that:
  1. (1)
    $K_{2N}\geq 0$,
  2. (2)
[…]
  3. (3)
    $\hat{K}_{2N}\to 1$ pointwise on $\mathbb{Z}^{d}$.
Next, we need that the mass of $K_{2N}$ is concentrated near the origin. But this can be proved by mimicking the proof of [11, Lemma 3.8] and we omit the details. Hence $\{K_{2N}\}$ is good kernel. This concludes the proof. ∎
Corollary 4.7.
Let $d\in\mathbb{N}$. Then $\mathcal{UCP}_{\mathcal{K}}(\mathcal{T}_{N}^{d}(PH))$ converges to $\mathcal{UCP}_{\mathcal{K}}(C(\Omega))$ in the Gromov-Hausdorff distance.
Remark 4.8.
As one more example of Gromov-Hausdorff convergence of CP maps, consider $V_{n}$, the $n$-dimensional irreducible representation of $SU(2)$. Let $S^{2}$ be the $2$-dimensional unit sphere. Then $\mathcal{UCP}_{\mathcal{K}}(L(V_{n}))$ converges to $\mathcal{UCP}_{\mathcal{K}}(C(S^{2}))$. This is a consequence of non-trivial results of van Suijlekom in section 3.3 of [16] and our Theorem 1. van Suijlekom produces the $C^{1}$-approximate order isomorphism and hence Theorem 1 can be applied.
Acknowledgement: The authors are thankful to Walter van Suijlekom, whose India visit under the Jubilee Chair Professorship of the Indian of Academy of Sciences was pivotal for this paper, as well as M. Leimbach for his comments. The first author is supported by a J C Bose Fellowship JCB/2021/000041 of SERB. The second author is supported by the Prime Minister’s Research Fellowship. The third author is supported by NBHM post-doctoral fellowship.
Declarations
On behalf of all authors, the corresponding author states that there is no conflict of interest.
This manuscript has no associated data.